It is claimed that a car is driven on the average less than 28700 kilometers per year with a population standard deviation of 2100 kilometers. To test this claim, a random sample of 58 car owners are asked to keep a record of the kilometers they travelled. Would you agree with this claim if the random sample showed an average of 22500 kilometers? Use a 0.01 level of significance


1
Expert's answer
2022-03-24T06:38:20-0400

"\mu=28700, \sigma=2100,n=58,\bar{x}=22500,\alpha=0.01."


Null and alternative hypotheses:

"H_0:\mu\geq28700;\\
H_1:\mu<28700."


Because "\sigma" is known and "n=58>30," we can use the z-test.

"z=\cfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\cfrac{22500-28700}{2100/\sqrt{58}}=-22.48."

In z-table, the area corresponding to "z=-22.48" is 0. Because the test is a left-tailed test, the P-value is equal to the area to the left of "z=-22.48," so, "P=0."

Because the P-value is less than "\alpha" =0.01, we reject the null hypothesis, there is enough evidence at the 1% level of significance to support the claim that the mean run is less than 28700 kilometers per year.